rizen is 5 years older than JB. Find their present ages if the product of Rizen's age 4 years from now and JB's age 5 years ago is 51
Answer by stanbon(57246)

rizen is 5 years older than JB. Find their present ages if the product of Rizen's age 4 years from now and JB's age 5 years ago is 51
--------------
NOW DATA:
JB is "x" yrs old ; Rizen is "x+5" yrs old:
-------------
4 yrs from NOW and 5 years from NOW DATA:
Jb will be "x+4" yrs old; Rizen will be "x+10" yrs old
-------------
EQUATION:
(x+4)(x+10) = 51
x^2+14x+40 = 51
x^2+14x-11 = 0
x=[-14+-sqrt(14^2-4*1*-11)]/2
x=[-14 +- sqrt(240)]/2
Positive answer:
x = 0.75 (JB's age NOW)
x+5 = 5.75 (Rizen's age NOW)
==============
Cheers,
stan H.